equation 4 = MNaOH = mKHP/(MWKHP(V2 ? V1)NaOH) Write the equation and determine all the partial derivatives that are used to calculate the uncertainty (error) in the NaOH concentration? Be sure to use eq(4). the question is number 8 on the pictures!

Experiment 3 -Preparing and Standardizing NaOH The goals are (1) to prepare and standardize a sodium hydroxide solution that is in a range of 0.095-0.11 M, (2) to perform a statistical analysis of the 3 trials, and (3) to determine the uncertainty in the final NaOH concentration value. Chemicals & Supplies 1. 2. 3. 4. solid NaOH oven dried solid potassium hydrogen phthalate (KHP) (Beury 211) 1 liter of boiled then cooled water (done previously) in a plastic bottle phenolphthalein indicator (50 mg in 50 mL EtOH+50 mL water) Solid sodium hydroxide (NaOH) is an inexpensive and readily available strong base. Both the solid and aqueous solutions of NaOH are extremely hygroscopic. As such, weighing this material must be done fairly quickly, and term storage because NaOH dissolves glass. Exposure of NaOH to air must be kept to a minimum because of its → NaHCO3). It is for this reason that boiled water must be used. The boiling process expels any dissolved CO2 storing the solid or an aqueous solution in a tightly capped bottle is necessary. Plastic bottles are required for long hygroscopic nature and its reaction with carbon dioxide to make the weak base, sodium bicarbonate (NaOH + CO2 Because of the unavoidable reactivity of NaOH with both water and CO, it is necessary to reference the concentration of our aqueous NaOH solution to another, less reactive acid solution. This process is known as standardization. We will be using potassium hydrogen phthalate (KHC3M,0s) as our primary standard, a high purity stable compound to which all of our values will be referenced. This crystalline monoprotic acid reacts with NaOH according to the reaction: NaOH + KHC8H404 → H2O + Na KC8H,04. To calculate the [NaOH), we use the following relationships (eq 1) (eq 2) (eq 3) (MV)NaOH = (MV)KHP MyaOH(V2-V1) NaOH = nKHP KHP MN mKHP (eq 4) NaOH It It is from eq(4) that the uncertainty (error) in the [NaOH] will be found. It is left as a pre-lab question for you to write the equation for the error in (NaOH] Be sure that you are not being wasteful with your NaOH solution because it will be used for 4 experiments. Be sure that you label your bottle with your name, contents and date. Store your bottle in the designated location for your lab section. Special safety considerations for NaOH are required, because it painlessly dissolves skin and hair. If you spill some on your skin, it will feel soapy. You are required to fill all burets below eye level to avoid splashing NaOH into your eyes. If you fail to fill your burets properly, you will be dismissed from lab for the day and receive a zero score.Numer 8 please, I dont need any of the other questions
Answer #1

Equation 4 is

MNaOH = mKHP / MWKHP (V2-V1)NaOH

so concentration of NaOH is a function of mass of KHP, mol wt of KHP and volume of NaOH

d(MNaOH) = f(d(mKHP), d(MWKHP), d((V2-V1)NaOH))

The total deviation of concentration can be calculated by taking partial derivatives of x with respect to each term

d(M_{NaOH}) = (\frac{\delta (M_{NaOH})}{\delta m_{KHP}})dm_{KHP}   , (\frac{\delta (M_{NaOH})}{\delta MW_{KHP}})dMW_{KHP} ,, (\frac{\delta (M_{NaOH})}{\delta (V2-V1))_{NaOH}})d(V2-V1))_{NaOH}

Now on squaring and crossing the terms and dividing by N-1 we will get the equation for propagation of error.

σ2(MNaOH )=(δMNaOH / δmKHP)2σ2mKHP + (δMNaOH / δMWKHP)2σ2MWKHP + (δMNaOH / δ(V2-V1)NaOH)2σ2V2-V1)NaOH